OpenStudy (anonymous):

Which of the following is equivalent to log32 4? 0.4 0.424 0.125 2.5

OpenStudy (anonymous):

@mathstudent55

jimthompson5910 (jim_thompson5910):

use the change of base law \[\Large \log_{32}(4) =\frac{\log(4)}{\log(32)} = ???\]

OpenStudy (anonymous):

0.4 :)